### v2

# From Models to Network Topologies: A Topology Inference Attack in Decentralized Federated Learning

Abstract
Federated Learning (FL) is widely recognized as a privacy-preserving machine learning paradigm due to its model-sharing mechanism that avoids direct data exchange. Nevertheless, model training leaves exploitable traces that can be used to infer sensitive information. In Decentralized FL (DFL), the topology, defining how participants are connected, plays a crucial role in shaping the model’s privacy, robustness, and convergence. However, the topology introduces an unexplored vulnerability: attackers can exploit it to infer participant relationships and launch targeted attacks. This work uncovers the hidden risks of DFL topologies by proposing a novel Topology Inference Attack that infers the topology solely from model behavior. A taxonomy of topology inference attacks is introduced, categorizing them by the attacker’s capabilities and knowledge. Practical attack strategies are designed for various scenarios, and experiments are conducted to identify key factors influencing attack success. The results demonstrate that analyzing only the model of each node can accurately infer the DFL topology, highlighting a critical privacy risk in DFL systems. These findings offer valuable insights for improving privacy preservation in DFL environments.
[…]
123
1 Introduction
Federated Learning (FL) has emerged as a novel framework for enabling privacy-preserving Machine Learning (ML), facilitating collaborative model training among distributed clients without the necessity of sharing raw data [25]. Conventional FL systems are predicated on a centralized architecture, referred to as Centralized FL (CFL), wherein a central entity is responsible for collecting, aggregating, and redistributing models to the clients. Nevertheless, this centralized architecture presents several challenges, such as processing bottlenecks and the single points of failure risk [3]. In response to these limitations, Decentralized FL (DFL) has been introduced, wherein model training and aggregation occur locally, and communication between nodes relies on a peer-to-peer (P2P) network. This approach eliminates dependency on a central server, mitigating risks associated with single points of failure [2].
Although FL, both centralized and decentralized, protect raw data privacy through its unique model-sharing mechanism, the training process inevitably leaves traces in the models. These traces can be exploited by malicious actors, potentially leading to inference attacks, including membership inference attacks, property inference attacks, and attribute inference attacks [20]. Previous studies have illustrated that such attacks pose a considerable risk to privacy within FL systems by elucidating sensitive information based on the behavior of the models [7]. While much of the existing work focuses on CFL, a notable research gap exists in exploring information leakage in DFL.
Overlay network topology defines how participants are interconnected in an FL system. In CFL, the client-server architecture enforces a fixed, star-shaped topology. However, DFL leverages the P2P network, enabling flexible node connections to form diverse topologies. Existing studies demonstrate that topology significantly influences DFL models’ robustness and privacy-preserving capabilities [10]. From a security perspective, exposure of the DFL network topology significantly facilitates malicious attacks. For example, in poisoning attack scenarios, topology awareness allows adversaries to pinpoint critical nodes, such as the central node in star topologies, thereby enabling them to amplify the poisoning intensity and enhance its propagation effectiveness [10]. Similarly, for network-level attacks, knowledge of key nodes within the DFL topology can lead to more targeted and effective Distributed Denial-of-Service (DDoS) attacks. Thus, topology information must be regarded as a critical asset, both in terms of optimizing DFL training performance and ensuring system security. However, theoretical analyses on how network topology affects model convergence remain sparse. Moreover, there is limited research investigating potential information leakage risks for this sensitive overlay topology data and strategies to safeguard this critical asset.
Therefore, this paper improves this research gap by proposing a novel overlay topology inference attack targeting DFL systems. The proposed attack leverages the behavioral traces generated by models to uncover sensitive overlay topologies. The main contributions of this work are as follows: (i) the formulation of a novel taxonomy of topology inference attacks, classifying these attacks based on the attacker’s capabilities and knowledge; (ii) the development of practical strategies tailored to various attack types, accompanied by a quantitative analysis of critical factors that influence the efficiency of these attacks; (iii) rigorous experimental evaluations conducted across a range of datasets and real-world network topology configurations to empirically validate the proposed attack strategies; and (iv) the provision of insights aimed at informing the design of effective defensive mechanisms to safeguard sensitive information and preserve privacy within DFL systems.
2 Background and Related Work
This section overviews inference attacks, encompassing their classifications and potential attack surfaces. Since there is a lack of research directly related to topology inference attacks on DFL, this paper examines topology inference studies conducted in other domains, including communication systems and social networks.
2.1 Inference Attacks
Inference attacks exploit ML models to extract sensitive information without direct access to the underlying data [29]. By analyzing model behaviors or outputs, adversaries can infer properties of the data, model parameters, or system structure. As FL grows in adoption for its privacy-preserving capabilities, it also faces increased vulnerability to such attacks [36]. Inference attacks vary in objectives and strategies, with common types including:
  • Membership Inference Attack: Determines whether a specific data sample was part of the training dataset, often using shadow models or prediction confidence scores to differentiate between training and unseen data [31]. This is critical in sensitive domains like healthcare or finance.
  • Model Inversion Attack: Reconstructs input data or infers its properties by iteratively optimizing inputs to match observed model outputs, posing risks in applications involving sensitive information like facial recognition or medical data [11].
  • Property Inference Attack: Infers characteristics of the training data, such as demographic distributions, by analyzing model updates or outputs. In FL, this can involve monitoring node-specific behaviors to deduce shared properties of local datasets [12].
  • Attribute Inference Attack: Targets specific attributes of training data, often by injecting malicious samples to influence the model’s learning process and make it vulnerable to data leakage [14].
These attacks not only compromise users’ sensitive information and undermine data security but also erode trust in FL’s privacy protection mechanisms. As a result, users may hesitate to participate in the training process, undermining the overall effectiveness of the FL system. However, existing research on inference attacks in FL primarily focuses on CFL, with limited exploration of privacy leakage in DFL. Moreover, these studies have not adequately addressed the critical role of DFL’s overlay topology.
2.2 Topology Inference
Topology inference is explored in various research domains, such as identifying source–destination paths in communication systems and uncovering interconnections among nodes in social networks.
2.2.1 Communication Systems
In communication systems, most protocols are designed to avoid ring formation, resulting in tree-like overlay topologies. Moreover, many topology inference studies commonly assume a single path between nodes [5]. A widely used approach involves leveraging ICMP-based protocols, such as the traceroute, to map network topologies effectively [21]. Tomography-based metrics, like packet loss rates, further aid in deducing a system’s topology: by observing loss rates from a source node to a target node, one can estimate the likelihood of a direct connection and thereby reconstruct the network [6]. Additionally, by measuring the timestamps of sent and received packets, it is possible to determine the number of forwarding hops, offering deeper insight into the overall network configuration [19].
2.2.2 Social Networks
Social networks generally refer to the social structure formed by individuals in society connected by a specific relationship. In social networks, user interactions (e.g., likes, retweets) serve as valuable signals for topology modeling and inference by capturing how information diffuses across the network [13]. However, this behavioral data is often incomplete, especially in large-scale settings, necessitating methods like noisy sparse subspace clustering to infer network structures from partial observations [33]. Meanwhile, social networks are dynamic and time-sensitive, and they frequently exhibit small-world properties.[15] incorporate these distributional patterns as prior knowledge in evolutionary models to better capture network changes over time.
Communication networks that often assume tree topologies and social networks are typically modeled as small-world networks; however, these assumptions do not hold for DFL network topologies. Moreover, common inference tools and metrics, such as traceroute, packet loss rate, or user interaction data, are not readily applicable in DFL contexts, creating the need for specialized inference metrics and strategies tailored to the particularities of DFL systems.
3 Overlay Topology in DFL
DFL is an ML paradigm that facilitates collaborative model training across multiple nodes without a central server. In DFL, each node updates its local model parameters using its private data and exchanges them with its directly connected neighbors. As the overlay topology in DFL significantly affects model convergence, this work presents the topology as an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ represents the edges indicating direct communication links between nodes. The adjacency matrix $A$ encodes the graph structure, where:
$A_{ij}=\begin{cases}1&\text{if nodes $i$ and $j$ are connected},\\ 0&\text{otherwise}.\end{cases}$ (1)
The degree matrix $D$ is a diagonal matrix where $D_{ii}=\text{degree}(i)+1$. The aggregation process is governed by the normalized aggregation matrix:
$P=D^{-1}(A+I),$ (2)
where $I$ is the identity matrix. This formulation ensures $P$ is row-stochastic ($\sum_{j}P_{ij}=1$), enabling consistent scaling during aggregation.
DFL proceeds iteratively over $T$ rounds, comprising local training and model aggregation steps. Let $M_{t}=[\theta_{1}^{(t)},\theta_{2}^{(t)},\dots,\theta_{N}^{(t)}]$ represent the models of all nodes at round $t$, where $\theta_{i}^{(t)}$ denotes the parameters of node $i$.
During the training stage, each node updates its model using its local dataset:
$M_{t}=\tilde{M}_{t-1}+\delta_{t},$ (3)
where $\delta_{t}$ represents local updates, and $\tilde{M}_{t-1}$ is the aggregated model from round $t-1$.
In the aggregation stage, nodes aggregate models from neighbors using the aggregation matrix $P$:
$\tilde{M}_{t}=PM_{t-1}.$ (4)
The model parameters after $T$ rounds can be expressed as:
$M_{T}=P^{T}M_{0}+\sum_{t=1}^{T}P^{T-t}\delta_{t}.$ (5)
Assuming $G$ is connected (i.e., no isolated nodes), the spectral radius of $P$ satisfies $\rho(P)<1$. Additionally, under bounded local updates (e.g., using SGD or Adam), there exists $C>0$ such that $\|\delta_{i}^{(t)}\|\leq C$ for all $t$ and $i$. As $T$ increases, the influence of the initial model diminishes:
$\|P^{T}M_{0}\|\leq\|P^{T}\|\|M_{0}\|\rightarrow 0\quad\text{as}\quad T\rightarrow\infty.$ (6)
The cumulative effect of local updates is:
$\sum_{t=1}^{T}P^{T-t}\delta_{t},$ (7)
where $\|P^{T-t}\delta_{t}\|$ is bounded by $\|P^{T-t}\|\|\delta_{t}\|$. Since $\|P^{T-t}\|\rightarrow 0$ as $T\rightarrow\infty$, the impact of individual updates diminishes, and the series converges due to $\rho(P)<1$.
Thus, it can be seen that the convergence behavior of DFL is influenced by the topology captured by $P$. Initially, the aggregation dynamics (powers of $P$) dominate, reducing the influence of $M_{0}$. Over time, local updates drive the learning process, modulated by the decaying powers of $P$. This interplay underscores the importance of overlay topology in determining convergence efficiency.
Given the critical role of overlay topology, it is essential to study potential information leakage through model updates and develop effective methods to protect sensitive topology information in DFL.
4 Problem Statement
Theoretical analysis of the DFL learning process reveals that overlay topology is crucial in determining system performance, robustness, and privacy. This section introduces the problem of topology inference and categorizes potential attack types based on the attacker’s knowledge and capabilities.
Consider a DFL system represented by an undirected graph $G=(V,E)$, where $V$ is the set of participating nodes, and $E$ is the set of edges representing direct communication links between nodes. The adjacency matrix $A$ of size $|V|\times|V|$ encodes the network topology, as defined in Equation (1).
An attacker aims to infer a predicted adjacency matrix $A^{\prime}$ that approximates the ground truth $A$ as closely as possible, thereby revealing sensitive topology information.
The effectiveness of topology inference attacks depends on the attacker’s knowledge and capabilities. The following assumptions are made. (i) Internal Adversary: The attacker is an internal participant in the DFL system and can identify a subset of nodes. (ii) Decoupled Information: Node-level information (e.g., models and datasets) and network-level information (e.g., connections) are treated as separate assets. For instance, an attacker may know a node’s model but not its connections.
Let $V^{\prime}\subset V$ represent a set of nodes known to the attacker and $E^{\prime}\subset E$ represent the subset of edges known to the attacker, where $\emptyset\subsetneq E^{\prime}\subsetneq E$ to represent partial knowledge. The attacker’s goal is to infer the entire edge set $E$ based on partial knowledge of $V^{\prime}$ and $E^{\prime}$.
For each node $i\in V$, the local model is represented as $M_{i}$, and the local dataset is represented as $D_{i}$. This paper assumes that models and data within a node are distinct assets. Thus, the attacker’s capabilities are defined as follows:
  • Model Knowledge: The attacker can access models of known nodes $V^{\prime}$, forming the set: $M^{\prime}=\{M_{i}\mid i\in V^{\prime}\}.$
  • Dataset Knowledge: The attacker may not have access to the datasets $D_{i}$ of these nodes. The datasets known to the attacker are denoted by: $D^{\prime}\subseteq\{D_{i}\mid i\in V^{\prime}\}\quad\text{or}\quad D^{\prime}=\emptyset.$
Based on the attacker’s knowledge and capabilities, the topology inference attack is classified into five scenarios:
  • Scenario 1: The attacker knows all node models and datasets, but only partial edge information ($M^{\prime}=M,D^{\prime}=D,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 2: The attacker knows all node models, but none of the datasets, and only partial edge information ($M^{\prime}=M,D^{\prime}=\emptyset,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 3: The attacker knows all node models and datasets, but does not have edge information ($M^{\prime}=M,D^{\prime}=D,E^{\prime}=\emptyset$).
  • Scenario 4: The attacker knows all node models, but neither datasets nor edge information ($M^{\prime}=M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
  • Scenario 5: The attacker knows partial models, but neither datasets nor edge information ($\emptyset\subsetneq M^{\prime}\subsetneq M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
For Scenario 1, since the attacker can control all nodes and knows partial information about the edges, it can be classified as a white-box attack. Scenarios 2, 3, and 4 can be categorized as gray-box attacks. For attack scenario 5, no information is available regarding the total number of nodes or edges, making it infeasible to fully reconstruct the network topology with the approaches considered. As a result, this work only considers the attack scenarios 1, 2, 3, and 4.
5 Topology Inference Attack
This section first explores the metrics that attackers can exploit to carry out topology inference attacks, and then designs attack strategies for various attack scenarios.
5.1 Attack Metrics
In a DFL system, attackers may leverage the local data and local models available at each known node. Additionally, DFL’s decentralized communication mechanism allows attackers to access models transmitted from neighboring nodes. Thus, this paper proposes six attack metrics that attackers can exploit:
  • Relative Loss. This metric evaluates how well a model trained on one node’s data generalizes to another node’s dataset. Consider a model $f_{i}$ trained on node $i$’s dataset $D_{i}$, and let $D_{j}$ be the dataset of another node $j$. The relative loss $\text{Relative Loss}_{i,j}$ measures the performance of $f_{i}$ on $D_{j}$. A lower relative loss indicates better transferability and generalization across nodes:
    $\text{Relative Loss}_{i,j}=\mathcal{L}(f_{i},D_{j})$ (8)
    where $\mathcal{L}(f,D)$ is the loss function used by model $f$ on dataset $D$.
  • Relative Entropy. This metric measures the uncertainty of $f_{i}$’s predictions on $D_{j}$. While relative loss focuses on correctness, relative entropy focuses on confidence. A model may be confident (low entropy) but still perform poorly if it is incorrect.
    $\text{Relative Entropy}_{i,j}=-\frac{1}{|D_{j}|}\sum_{x\in D_{j}}\sum_{k}y_{k}(x)\log(f_{i,k}(x))$ (9)
    where $y_{k}(x)$ is the true label distribution for sample $x$ and $f_{i,k}(x)$ is the predicted probability of class $k$ under model $f_{i}$.
  • Relative Sensitivity. This metric quantifies the sensitivity of model $f_{i}$’s predictions on dataset $D_{j}$ using the Jacobian norm. Relative sensitivity could capture the model’s susceptibility to small input perturbation.
    $\text{Relative Sensitivity}_{i,j}=\frac{1}{|D_{j}|}\sum_{x\in D_{j}}\left\|\frac{\partial f_{i}(x)}{\partial x}\right\|_{F}$ (10)
    where $\frac{\partial f_{i}(x)}{\partial x}$ is the Jacobian of $f_{i}$’s outputs with respect to input $x$, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
  • Cosine Similarity. This metric compares the direction of two parameter vectors $\mathbf{a}$ and $\mathbf{b}$ derived from different models. If both vectors point in similar directions, their cosine similarity is high. In DFL, a high cosine similarity suggests that the two nodes frequently aggregate their parameters, resulting in more closely aligned models:
    Cosine $\text{Cosine Similarity}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\,\|\mathbf{b}\|}$ (11)
    where $\mathbf{a}\cdot\mathbf{b}$ is the dot product, and $\|\mathbf{a}\|$, $\|\mathbf{b}\|$ are the L2 norms of the vectors $\mathbf{a}$ and $\mathbf{b}$, respectively.
  • Euclidean Similarity. It quantifies how similar two vectors are based on the Euclidean distance between them. A higher Euclidean similarity indicates that the two models are numerically closer, while a smaller distance suggests they differ significantly:
    Euclidean $\text{Euclidean Similarity}(\mathbf{a},\mathbf{b})=\frac{1}{1+\sqrt{\sum_{i=1}^{n}(a_{i}-b_{i})^{2}}}$ (12)
    where $a_{i}$ and $b_{i}$ are the $i$-th elements of $\mathbf{a}$ and $\mathbf{b}$.
  • Curvature Divergence. This metric quantifies model divergence by analyzing implicit second-order information from consecutive parameter updates, reflecting curvature characteristics without explicit data dependence. Specifically, it compares parameter update vectors to capture similarity in model evolution indirectly:
    Curvature $\text{Curvature Divergence}(\Delta\mathbf{a},\Delta\mathbf{b})=\frac{\|\Delta\mathbf{a}-\Delta\mathbf{b}\|_{2}}{\frac{1}{2}(\|\Delta\mathbf{a}\|_{2}+\|\Delta\mathbf{b}\|_{2})}$ (13)
    where $\Delta\mathbf{a}=\mathbf{a}^{t}-\mathbf{a}^{t-1}$ and $\Delta\mathbf{b}=\mathbf{b}^{t}-\mathbf{b}^{t-1}$ denote parameter updates between rounds $t-1$ and $t$. A lower value indicates higher curvature divergence.
Relative metrics, including relative loss, entropy, and sensitivity, require access to both local models and local datasets, thus making them suitable for attack scenarios 1 and 3. While similarity metrics, i.e. cosine, Euclidean, and curvature similarity, depend only on the models and do not require dataset access, they are appropriate for attack scenarios 2 and 4.
5.2 Attack Strategies
In a topology inference attack, the attacker’s objective is to construct a predicted adjacency matrix $A^{\prime}$ that closely approximates the ground-truth adjacency matrix $A$.
5.2.1 Supervised Attack Strategy.
In attack scenarios 1 and 2, where the attacker has access to partial ground-truth edge labels, this work introduces EDGEPRE, a supervised edge classification framework for topology reconstruction. As shown in Algorithm 1, the model learns to infer the full adjacency matrix from node-level behavioral features and a small set of labeled node pairs.
Model Architecture.
Given $N$ nodes and their behavioral metrics $X\in\mathbb{R}^{N\times d}$, EDGEPRE directly uses an MLP-based edge classifier without any message passing. For each labeled pair $(i,j)$, the concatenated feature pair $[x_{i}\|x_{j}]$ is passed to a multilayer perceptron (MLP) decoder that outputs the predicted edge probability:
$\hat{y}_{ij}=\sigma(f_{\text{dec}}([x_{i}\|x_{j}])).$
Optionally, feature interaction terms such as $x_{i}*x_{j}$ and $|x_{i}-x_{j}|$ can be included to enhance the representation.
Training Objective.
The model is optimized using binary cross-entropy over the labeled edge set:
$\mathcal{L}=\sum_{(i,j,y_{ij})\in\mathcal{P}}\text{BCE}(\hat{y}_{ij},y_{ij}).$
Adjacency Reconstruction.
After training, EDGEPRE generalizes to all node pairs and produces a reconstructed soft adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$, where each entry is computed via:
$A^{\prime}_{ij}=\sigma(f_{\text{dec}}([x_{i}\|x_{j}])).$
Algorithm 1 EDGEPRE (MLP-based)
1:Node features $X\in\mathbb{R}^{N\times d}$, labeled edge set $\mathcal{P}_{\text{train}}=\{(i,j,y_{ij})\}$, number of epochs $T$
2:Reconstructed adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$
3:Initialize MLP decoder $f_{\text{dec}}$
4:for $t=1$ to $T$ do
5:     for each labeled pair $(i,j,y_{ij})\in\mathcal{P}_{\text{train}}$ do
6:         $h_{ij}\leftarrow[x_{i}\|x_{j}]$
7:         $\hat{y}_{ij}\leftarrow\sigma(f_{\text{dec}}(h_{ij}))$
8:     Compute loss: $\mathcal{L}\leftarrow\sum_{(i,j)}\text{BCE}(\hat{y}_{ij},y_{ij})$
9:     Update decoder parameters via gradient descent
10:Inference phase
11:for all $(i,j)$ where $1\leq i,j\leq N$ do
12:     $h_{ij}\leftarrow[x_{i}\|x_{j}]$
13:     $A^{\prime}_{ij}\leftarrow\sigma(f_{\text{dec}}(h_{ij}))$
14:return $A^{\prime}$
5.2.2 Unsupervised Attack Strategy.
For attack scenarios 3 and 4, where no edge labels are available, this work proposes INFERGAT, an unsupervised structure inference method based on a GAT encoder-decoder framework. The model predicts the latent adjacency purely from node-level behavioral metrics without requiring ground-truth edges.
Model Architecture.
The input $X\in\mathbb{R}^{N\times d}$ contains attack-related features such as similarity or entropy. A GAT encoder outputs embeddings $Z=f_{\text{enc}}(X)$, and a symmetric MLP decoder estimates pairwise edge strength via $A^{\prime}_{ij}=\sigma(\text{MLP}([z_{i}\|z_{j}]))$.
Training Objective.
The model is trained to reconstruct the behavioral structure with mean squared error:
$\mathcal{L}=\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$ (14)
The final output is the predicted adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$.
Algorithm 2 INFERGAT
1:Attack metrics $X\in\mathbb{R}^{N\times d}$, learning rate $\eta$, epochs $T$
2:Predicted adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$
3:Initialize GAT encoder $f_{\text{enc}}$ and decoder $f_{\text{dec}}$ with parameters $\theta$
4:for $t=1$ to $T$ do
5:     $Z\leftarrow f_{\text{enc}}(X)$triangleright Node embeddings via multi-head GAT
6:     $A^{\prime}_{ij}\leftarrow f_{\text{dec}}(z_{i},z_{j}),\quad\forall i,j$triangleright Decoder predicts edge weights
7:     $A^{\prime}\leftarrow\frac{1}{2}(A^{\prime}+A^{\prime\top})$triangleright Symmetrize adjacency
8:     $\mathcal{L}\leftarrow\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$triangleright Reconstruction loss
9:     $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$triangleright Gradient descent update
10:return $A^{\prime}$
EDGEPRE and INFERGAT represent two attack strategies for topology inference in DFL. EDGEPRE leverages partial edge supervision to perform explicit edge classification, while INFERGAT relies solely on behavioral metrics to reconstruct structural patterns in a self-supervised manner. Together, they cover a broad range of adversarial capabilities and reflect realistic threat models across different attack scenarios.
6 Evaluation
This work empirically validates the effectiveness of the proposed topology inference attack. First, it identifies which metrics yield stronger attack performance. Building on these results, it then thoroughly evaluates the proposed attack strategies under various conditions, including different datasets, topologies, node counts, and data preprocessing methods. This comprehensive assessment demonstrates the robustness and applicability of the proposed topology inference attack.
6.1 Experimental Setups
This section describes the configurations employed in the experiments.
A. Dataset and Model: All experiments are conducted on five datasets: MNIST, Fashion-MNIST, CIFAR10, SVHM, and ImageNet100, widely used in FL and inference attack benchmarks.
  • MNIST [23]: Contains 60,000 grayscale training images and 10,000 test images of handwritten digits (0–9) at 28$\times$28 resolution. A two-hidden-layer MLP (256 and 128 neurons) trained with Adam (lr=1e-3) is used.
  • FMNIST [35]: Similar size and format to MNIST, but with 10 categories of clothing items. A CNN with two convolutional layers (32 and 64 filters, kernel size 3$\times$3) is employed.
  • CIFAR10 [22]: Contains 60,000 32$\times$32 RGB images from 10 categories. The training set is only treated with basic image normalization. The image pixel values are normalized to keep the original image content unchanged. A MobileNet [30] is used for training.
  • SVHN [26]: Composes of over 600,000 32$\times$32 RGB images of house numbers collected from Google Street View. Each image contains a centered digit from 0 to 9. A ResNet9 [17] is adopted as the training model.
  • ImageNet10: A curated subset of the ImageNet [8], containing 10 classes selected from the original ImageNet hierarchy. This version [1] includes approximately 15,000 images with varied object categories and balanced class distributions. A lightweight PoolFormer-S12 [37] model is used for training.
B. DFL Topology Setting: All experiments are conducted on DFL systems comprising two categories of topologies: synthetic and real-world. The synthetic topologies include ring, star, and three Erdős-Rényi (ER) graphs with connection probabilities $p=$0.3, 0.5, and 0.7 [9]. The real-world topologies consist of 29 network topologies collected from the SNDlib [27] and DEFO [16] repositories, which reflect realistic communication and infrastructure networks. Detailed descriptions and properties of these topologies are provided in the supplementary material.
C. Federation Setting: All experiments employ the following DFL configurations:
  • Number of Nodes: For synthetic topologies, the number of nodes is set to 10, 20, and 30. For real-world topologies, the number of nodes ranges from 12 to 161, depending on the specific topology.
  • Total Rounds: The total number of communication rounds is determined by the rule rounds = number of nodes. The mechanism ensures efficient inter-node communication, thereby enabling consistent scalability of training time as the network size increases.
  • Local Epochs: Two scenarios (3 and 10 local epochs) are considered to evaluate how varying levels of local overfitting affect the attack’s effectiveness.
  • Data Distribution: Both Independent and Identically Distributed (IID) and non-IID $\alpha=0.1$ data distributions are considered to evaluate the impact of data heterogeneity.
D. Attack Models: To evaluate the effectiveness of the proposed methods, EDGEPRE is compared against classical supervised classifiers, including Logistic Regression [34], Support Vector Machine (SVM) [18], and Random Forest (RF) [4]. For the unsupervised attack strategy, INFERGAT is evaluated against standard clustering-based methods, including K-Means [24], Gaussian Mixture Model (GMM) [28], and Spectral Clustering [32].
E. Evaluation Metric: Since the topology inference attack is formulated as a binary classification problem, the F1-Score set is used as the evaluation metric.
6.2 Selection of Attack Metrics
Figure 1: Normalized distributions of attack metrics in a 20-node star-topology DFL system trained on the FMNIST dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection.
Selecting appropriate attack metrics is critical to the success of inference attacks. An effective attack metric should distinguish between connected and non-connected nodes as clearly as possible. To this end, the first part of this experiment examines the effectiveness of the six proposed attack metrics by analyzing their respective distributions.
Figure 1 illustrates the distributions of these metrics for a 20-node star topology DFL system trained on the FMNIST dataset, comparing directly connected nodes (edge group) with those not (non-edge group). To facilitate direct comparison, values have been normalized.
The results indicate that the relative loss for connected nodes is notably lower than that for non-connected nodes, with minimal overlap between the respective distributions. By contrast, the relative entropy and sensitivity metrics exhibit substantially greater overlap. Since computing the relative loss requires access to local models and data, it has been employed in attack scenarios 1 and 3.
Regarding similarity-based metrics, the cosine similarity between connected models is distinctly higher than that between non-connected models, displaying no overlap between the distributions. On the other hand, the curvature divergence and Euclidean similarity show a considerably more significant degree of overlap. Thus, the cosine similarity has been adopted as the attack metric in scenarios where only local models are available (attack scenarios 2 and 4). The complete set of experimental results for all datasets and topologies is in the Appendix.
6.3 Attack Performance
After establishing the attack metrics, the proposed topology inference attacks were evaluated on five datasets, encompassing four attack scenarios and different topologies. The results in Table 1 demonstrate that the proposed methods, EDGEPRE and INFERGAT, are highly effective, achieving the best performance across all four attack scenarios and five datasets.
Table 1: Averaged Topology inference performance (F1-Score) under four attack scenarios across five datasets.
Scenario Algorithm CIFAR10 FMNIST ImageNet10 MNIST SVHN
SC1 Logistic 0.704 0.773 0.752 0.795 0.788
RF 0.844 0.845 0.879 0.844 0.851
SVM 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.851 0.875 0.840 0.876 0.874
SC2 Logistic 0.704 0.771 0.752 0.794 0.787
RF 0.844 0.845 0.879 0.844 0.851
SVC 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.850 0.875 0.843 0.876 0.872
SC3 GMM 0.723 0.443 0.482 0.669 0.679
Kmeans 0.727 0.715 0.618 0.724 0.734
Spectral 0.597 0.638 0.591 0.607 0.648
INFERGAT 0.800 0.788 0.733 0.799 0.802
SC4 GMM 0.576 0.662 0.685 0.670 0.688
Kmeans 0.730 0.776 0.741 0.785 0.795
Spectral 0.683 0.645 0.636 0.637 0.627
INFERGAT 0.797 0.810 0.743 0.807 0.812
It can also be observed that from SC1 to SC4, the attacker’s accessible knowledge gradually decreases, leading to a corresponding decline in attack performance. In SC1, the inferred topologies achieve an F-score exceeding 85% on all five datasets, whereas in SC4, the score drops to approximately 80%. This indicates that prior knowledge has a significant impact on the success rate of topology inference. Comparing SC1 and SC2, the difference in performance is relatively small, suggesting that combining data and model knowledge offers limited additional benefit over model-only knowledge. This implies that the model itself already encodes sufficient information to support effective topology inference.
A. Effect of Topology Density. To investigate the influence of topological characteristics on topology inference attacks, Figure 2 presents the F1-Scores achieved by each attack scenario under five in the CIFAR10 dataset. Topology density represents the ratio of the actual number of edges to the maximum possible number of edges, calculated as $\frac{2E}{N(N-1)}$. This metric reflects the overall connectivity or sparsity of a graph.
As shown in Figure 2, across four different attack scenarios, a consistent trend was observed: as the graph’s density increases, the attack success rate of the inference models (EDGEPRE and INFERGAT) tends to decrease. This is primarily because denser topologies lead to increased information fusion, resulting in higher similarity among local models. In highly dense networks, the models become increasingly homogeneous, making it more difficult for the inference algorithms to distinguish whether a connection exists between nodes. In contrast, sparse topologies exhibit greater heterogeneity among nodes despite being limited by data sparsity and class imbalance. This heterogeneity simplifies the inference task, allowing the attack models to make more accurate predictions.
Figure 2: Impact of topology density on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
B. Effect of Network Size. To evaluate the scalability of the proposed inference algorithms under different attack scenarios, Figure 3 illustrates the performance variation across four attack scenarios as the network size increases. The results show a consistent upward trend in attack effectiveness with increasing numbers of nodes, demonstrating the strong scalability of the proposed methods. On one hand, the algorithms accurately infer the underlying topology of DFL networks across varying levels of density, indicating their robustness in handling data imbalance challenges. On the other hand, even in large-scale networks with up to 161 nodes, the inference models maintain strong and stable performance, further validating the stability and scalability of the proposed topology inference approaches.
Figure 3: Impact of network size on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
7 Mitigation
The effectiveness of the proposed topology inference attacks indicates that topology in DFL systems can be discerned solely by analyzing model behaviors. This observation highlights the security and privacy concerns of DFL systems. This section proposes defense strategies to mitigate these vulnerabilities within DFL environments.
Figure 4: Impact of local training epochs, data augmentation, and data heterogeneity on the F1-Score of topology inference attacks in the CIFAR10 dataset.
7.1 Reduce Overfitting
As illustrated in Figure 4, each bar represents a specific training configuration: gray bars denote IID data distributions without augmentation, blue bars correspond to non-IID data distributions (Dirichlet $\alpha=0.1$), and orange bars indicate the use of data augmentation under IID settings.
Consistent with findings in prior inference attack literature, the success of topology inference attacks is closely tied to the degree of overfitting in the target model. To validate this, experiments were conducted on the CIFAR-10 dataset under four attack scenarios. By increasing the number of local training epochs from 3 to 10, the level of overfitting was elevated. Results show that, across all examined topologies, this change consistently improves the F1-Scores of attacks. This observation confirms that overfitting exacerbates the risk of topological information leakage in DFL systems.
To counteract overfitting, data augmentation techniques were employed, specifically random cropping with 4-pixel padding. The impact is evident in the substantial performance decline observed for the orange bars across all scenarios and topologies. This demonstrates that data augmentation not only enhances the generalization ability of local models but also effectively mitigates the success of topology inference attacks.
7.2 Data Heterogeneity
Furthermore, data distribution among clients plays a critical role in attack effectiveness. Non-IID data were simulated using a Dirichlet distribution with $\alpha=0.1$ to assess this. The results, represented by the blue bars, indicate a noticeable decline in attack performance across all scenarios, with the effect being most prominent in scenarios 3 and 4. This suggests that increased data heterogeneity introduces higher model divergence, which in turn hampers the inference model’s ability to exploit overfitted representations for edge prediction.
Overall, these findings imply that reducing overfitting through fewer local training epochs or data augmentation and increasing data heterogeneity are promising directions for mitigating topology inference risks in decentralized federated learning.
8 Conclusion and Future Work
This work introduces a novel topology inference attack against DFL, exposing critical vulnerabilities related to privacy and information leakage. By analyzing local models, attackers can accurately infer the overlay topology, one of DFL’s most sensitive assets, highlighting the system’s susceptibility to privacy breaches. The study explores various attack scenarios and develops tailored metrics, models, and algorithms, with experiments confirming the feasibility and effectiveness of these attacks. Furthermore, network size, density, model overfitting, and data heterogeneity significantly influence attack success. Mitigation strategies, such as data augmentation and increasing data heterogeneity, can enhance model generalization and reduce the risk of data leakage.
The present work focuses on attack scenarios 1 through 4 and does not yet address the most challenging attack scenario 5. Future research plans include exploring feasible strategies for Scenario 5 and extending the evaluations to a wider range of datasets and topologies. Besides, more attack metrics and strategies will be developed to improve attack effectiveness.